1.       Solve the exact differential equation (𝗑 − 2𝑒𝑦)𝑑𝑦 + (𝑦 + 𝗑 sin 𝗑)𝑑𝗑 = 0.

2.     Solve the differential equation 𝗑 𝑑𝑦 + 2𝑦 = 𝗑4.

1
Expert's answer
2022-02-04T10:08:38-0500

1.


"P(x,y)=y+x\sin x, \dfrac{\partial P}{\partial y}=1"

"Q(x,y)=x-2e^y, \dfrac{\partial Q}{\partial x}=1"

"\dfrac{\partial P}{\partial y}=1=\dfrac{\partial Q}{\partial x}"

Write the system of two differential equations that define the function "u(x,y)"


"\dfrac{\partial u}{\partial x}=y+x\sin x"

"\dfrac{\partial u}{\partial y}=x-2e^y"

"u=\int(x-2e^y)dy+\varphi(x)"

"u=xy-2e^y+\varphi(x)"


"\dfrac{\partial u}{\partial x}=y+\varphi'(x)=y+x\sin x"

"\varphi'(x)=x\sin x"


"\varphi(x)=\int x\sin xdx"

"\varphi(x)=-x\cos x+\sin x+C"


"u=xy-2e^y-x\cos x+\sin x+C"

"-xy+2e^y+x\cos x-\sin x=C"

2.


"x\dfrac{dy}{dx}+2y=x^4"

"y'+\dfrac{2}{x}y=x^3"

Integration factor


"\mu(x)=e^{\int(2/x)dx}=x^2"

"x^2y'+2xy=x^5"

"d(x^2y)=x^5 dx"

Integrate


"\int d(x^2y)=\int x^5 dx"

"x^2y=\dfrac{x^6}{6}+C"

"y=\dfrac{x^4}{6}+\dfrac{C}{x^2}"